OpenStudy (anonymous):

If x and y are non-zero numbers such that (6x+5y)/(3y-2x)=7, then x:y =?

OpenStudy (anonymous):

start with \[6x+5y=7(3y-2x)\] then you can solve for \(x\) in terms of \(y\)

OpenStudy (anonymous):

4:5 ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Re-arrange:\[\bf 6x+5y=7(3y-2x) \implies 6x+5y=21y-14x\]Now solve for x/y:\[\bf \implies 20x=16y\]\[\bf \implies \frac{ x }{ y }=\frac{ 16 }{ 20 }=\frac{ 4 }{ 5 }=4:5\]This is pretty much the same thing as phi's approach to the last question you posted. Re-arrange and solve for x/y. @kryton1212